Draw 11 rectangles with a perimeter of 24cm. For the width,use natural numbers, starting at 1.
Found 2 solutions by Alan3354, math_helper:
Answer by Alan3354(69443)
Answer by math_helper(2461)

You should try this yourself. You won't learn anything if someone else does this task for you.
Perimeter is twice the length + twice the width
P = 2L + 2W —> 2L = P-2W —> L=(1/2)(P-2W)
So if the first one has W=1, then 2L= 24-2(1) —> 2L = 22 —> L = 11
Your first rectangle has W=1 and L=11.

Set W=2, 2L = 24-2(2) = 20 —> L=10
etc.